OpenStudy (ebonyserrano):

On Saturday, a local hamburger shop sold a combined total of 327 hamburgers and cheeseburgers. The number of cheeseburgers sold was two times the number of hamburgers sold. How many hamburgers were sold on Saturday?

OpenStudy (ebonyserrano):

@ShadowLegendX

OpenStudy (shadowlegendx):

All you have to do is follow the same steps I showed you in the last problem ;o

OpenStudy (seratul):

X=Hamburgers 2X=Cheeseburgers 2x+x=327 Solve for x. Then replace x.

OpenStudy (shadowlegendx):

Let x = # of hamburgers sold Let y = # of cheeseburgers sold Since there are only these two types of burgers being sold, we can infer that, x + y = 327 But with this piece of information: "The number of cheeseburgers sold was two times the number of hamburgers sold." We can infer that y = 2x So going back to our first equation, we can replace y with 2x x + 2x = 327 Like terms 3x = 327 Divide the entire equation by 3 x = 109 109 hamburgers were sold on Saturday